# Higgs Inflation Model with Small Non-Minimal Coupling Constant

### v4

@@ 2.3 Vacuum of the Higgs field in the context of cosmology
The dependence of $U_{eff}^{(tree)}(\phi,\zeta;{\mathcal{M}})$, eq.(2.16), on $\zeta$, which, in turn, changes during the entire cosmological evolution (from inflation to transition to vacuum), plays a decisive role in all the physical results of the theory. Below, the TMT-effective potential will be obtained using eqs.(2.16) and (2.17). From now on we will asume that $V_{1}<0$, $V_{2}<0$ and $(electroweak\,scale)^{4}\lll|V_{1}|\sim|V_{2}|<M_{P}^{4}$ and we will see that by choosing the integration constant $\mathcal{M}$ one can ensure that the TMT-effective potential has a minimum at some value $\phi|_{min}=\sigma\lll M_{P}$. One can treat $\sigma$ as a vacuum expectation value (VEV) of $\phi$. The corresponding value $\zeta_{v}$ in the vacuum is obtained by making use of the constraint (2.17)
@@
Then eq.(2.19) reduces to an equation with a canonically normalised scalar field
@@
Here we neglected $\frac{\xi\sigma^{2}}{M_{P}^{2}}\lll 1$ in $\Omega$. One should pay special attention that in order to obtain equation (2.21) in canonical form, it was necessary to devide eq.(2.19), considered near the vacuum, by a constant $\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}$. In secs.2.4 and 6.1 we will return to the discussion of the source of this question in more general case and will see that it is directly related to the K-essence type structure of the theory, which is responsible for the important results of the paper.
@@
where we introduced the notation $\Lambda({\mathcal{M}})$ to emphasize that with an appropriate choice of ${\mathcal{M}}$, models with a very wide range of values of the cosmological constant can be realized. The natural constraint on $\Lambda$ is $0\leq\Lambda\lesssim(electroweak\,scale)^{4}$ and it is achieved if ${\mathcal{M}}^{2}\approx 4V_{1}V_{2}$. With the choice of the integration constant [^5] By an appropriate choice of $\delta$ it is possible to ensure that $\Lambda=0$. However, this is nothing more than fine tuning, that is, in this way we do not solve the CC problem.
@@
$\sigma^{2}=\frac{\zeta_{v}-b_{p}}{b_{p}+\zeta_{v}}\cdot\frac{m^{2}}{\lambda}\left[1+\mathcal{O}\left(\xi\frac{\sigma^{2}}{M_{P}^{2}}\right)\right].$ (2.28)
@@
and the notations [^6] We use the subscript $\sigma$ here to explicitly indicate that we are dealing with vacuum values of the TMT-effective parameters. Indeed, one might expect that, as follows from (2.19) and the constraint (2.17), $\phi$-dependent TMT-effective parameters $\lambda_{eff}$ and $m^{2}_{eff}$ appear as $\phi$ moves away from the vacuum. However, this is only partly true, namely for $\sigma<\phi\lll M_{P}$. As we will see in sections 2.4 and 4, substituting $\zeta$ from the constraint (2.17) into eq.(2.16) makes the TMT-effective potential essentially non-polynomial for $\phi\gtrsim 10^{-4}M_{P}$, which does not allow us to use the terms TMT-effective parameters $\lambda_{eff}$ and $m^{2}_{eff}$ in the general case.
@@
are introduced for the values of quartic selfinteraction and squared mass parameters near the vacuum $\sigma$. Note that up to corrections of the order $\sim\mathcal{O}\bigl{(}\xi\frac{\sigma^{2}}{M_{P}^{2}}\bigr{)}$
$\sigma^{2}=\frac{\zeta_{v}-b_{p}}{b_{p}+\zeta_{v}}\cdot\frac{m^{2}}{\lambda}=\frac{m_{\sigma}^{2}}{\lambda_{\sigma}}.$ (2.31)
If we restrict ourselves to studying the TMSM near vacuum, then the parameter $m$ will remain undefined, as well as the mass parameter in the GWS theory. The presence in the relations (2.30) of factors depending on the model parameters $\zeta_{v}=\sqrt{\frac{V_{1}}{V_{2}}}$, $b_{k}$ and $b_{p}$ means that there is a very wide range of these parameters in which the required VEV SM can be ensured [^7] However, in the model studied in this paper, we completely excluded gauge bosons and fermions and, accordingly, completely ignored the need to ensure gauge invariance. A suitably extended model that modifies the above results and allows adequate comparison with the GWS theory is studied in ref. [22], where it is shown that the VEV of the Higgs field, which should be compared with the VEV in the GWS theory, is not the value of $\sigma$ given by eq.(2.31), but the value of $\sqrt{\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}}\sigma$. As will be shown in ref.[22], demonstrating this requires taking into account the need to distinguish between the description of SM in the so-called cosmological frame (CF) and in the local particle physics frame (LPPF). From this point of view, in this paper we restrict ourselves to the work in the CF.. The fundamentally new effect of the model proposed in this paper is that, thanks to these three model parameters, there is also freedom in the choice of the parameter $\lambda$. As we will see in section 3, this allows the implementation of the Higgs inflation model with a small constant $\xi$ of the non-minimal Higgs coupling to the scalar curvature.
@@ 2.4 General view of the TMSM
In section 2.2, TMT-effective equations in the Einstein frame were obtained. To understand the main features of TMSM at energies above the electroweak scale, it is necessary to return to these equations, taking into account the results of section 2.3 relating to what happens near the Higgs vacuum in the CF (see footnote 5). First, from the requirement that the TMT-effective cosmological constant be negligibly small, we obtained that, neglecting insignificant corrections, the integration constant is $\mathcal{M}=2\sqrt{V_{1}V_{2}}$, eq.(2.24). Secondly, the scalar $\zeta$ in vacuum is equal to $\zeta_{v}=\sqrt{\frac{V_{1}}{V_{2}}}$, eq.(2.25). Therefore the term $\mathcal{M}+2\zeta V_{2}$ in $\phi$-equation (2.19) must be replaced by $2|V_{2}|(\zeta_{v}-\zeta)$.
$\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\Bigg{(}\frac{b_{k}-\zeta}{(1+\zeta)\Omega}\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}\phi\Bigg{)}+\frac{b_{p}+\zeta}{(1+\zeta)^{2}\Omega^{2}}\lambda\phi^{3}+\frac{b_{p}-\zeta}{(1+\zeta)^{2}\Omega^{2}}m^{2}\phi$
$+\frac{2\xi\phi}{(1+\zeta)M_{P}^{2}\Omega^{3}}\left[-2|V_{2}|(\zeta_{v}-\zeta)-\frac{\lambda}{4}\phi^{4}+\frac{m^{2}}{2}\phi^{2}-(1+\zeta)\Omega X_{\phi}\right]=0.$ (2.32)
Note that in the vacuum the $2|V_{2}|(\zeta_{v}-\zeta)$ term is zero.
From now on, we will use the following parametrization for $V_{2}$
@@
After substituting $\zeta(\phi,X_{\phi})$, eq.(2.17), into the $\zeta$-dependent form of $T_{\mu\nu}^{(eff)}$ in eqs.(2.15) and (2.16) with subsequent simple but rather lengthy algebraic calculations, the expression for the TMT-effective energy-momentum tensor reduces to
@@ 3.1 About the shape of the TMT-effective potential and other features of the model
The model contains 6 parameters $V_{1}$, $V_{2}$, $b_{p}$, $b_{k}$, $\lambda$ and $m$. Instead of $V_{2}<0$ it is convenient to use parameter $q^{4}\leq 1$ defined by the relation (2.33). Instead of $V_{1}<0$ we use $\zeta_{v}$ defined by eq.(2.25). Further in this paper, for simplicity, we choose
$V_{1}=V_{2}\quad\Longleftrightarrow\quad\zeta_{v}=1.$ (3.1)
It will be clear from what follows that this choice is made without loss of generality. As a specific example of small $\xi$ from now on, we will always use $\xi=\frac{1}{6}$. We also choose
@@
However, as can be seen from eq.(2.32), the redefinition (3.3) absorbs only the factor $\Omega^{-1}$, but does not provide canonical normalization of the field $\varphi$. The reason is that the TMT-effective action (2.38), (2.39) has the structure of K-essence type models. Thus, we are limited to only the redefinition (3.3) and arrive at the TMT-effective action in terms of $\varphi$
@@
The last expression for ${\tilde{K}}_{2}(\varphi)$ takes into account that $X_{\phi}=\cosh^{2}z\cdot X_{\varphi}$. The TMT-effective energy-momentum tensor given by eq.(2.34) now takes the following form
$T_{\mu\nu}^{(eff)}=K_{1}(\varphi)\cdot\big{(}\varphi_{,\mu}\varphi_{,\nu}-\tilde{g}_{\mu\nu}X_{\varphi}\big{)}-{\tilde{K}}_{2}(\varphi)\cdot\frac{X_{\varphi}}{M_{P}^{4}}\Big{(}\varphi_{,\mu}\varphi_{,\nu}-\frac{1}{2}\tilde{g}_{\mu\nu}X_{\phi}\Big{)}+\tilde{g}_{\mu\nu}U_{eff}^{(tree)}(\varphi).$ (3.11)
@@
Figure 2: The curve labeled 1 is a plot of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, defined by eqs.(3.6) and (3.7), has one plateau due to the choice of $q^{4}=3\cdot 10^{-10}$, i.e. $|V_{1}|=|V_{2}|\approx(10^{16}GeV)^{4}$, while other model parameters are the same as in Fig. 1. The curves labeled 2 and 3 are plots of the scalar function $\zeta(\varphi)$ defined by the constraint (3.1), where the terms $\propto\frac{X_{\varphi}}{M_{P}^{4}}$ are assumed to be negligible compared to the other terms. Curve 2 is the plot of $\zeta(\varphi)$ if the model parameter $b_{p}=0.55$ is chosen, and the plot intersects $\zeta=0$ at $\varphi=4.25M_{P}$. Curve 3 is the plot of $\zeta(\varphi)$ if the model parameter $b_{p}=0.5(1+10^{-8})$ is chosen, and the plot intersects $\zeta=0$ at $\varphi\approx 14M_{P}$.
@@
    Let us now analyze in more detail the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$ with one plateau. With the choice described by eqs.(3.1) and (3.2), and using the parameter values $\lambda=2.3\cdot 10^{-11}$,   $\xi=\frac{1}{6}$,   $m=0.7GeV$, one can see that the TMT-effective potential with one plateau can be achieved [^9] For $\varphi>\frac{M_{P}}{\sqrt{\xi}}=\sqrt{6}M_{P}$, we find that $F(\varphi)\approx\frac{1}{2}$ with a relative error $\frac{|F(\varphi)-\frac{1}{2}|}{1/2}<2.2\cdot 10^{-3}$. if $q^{4}\lesssim 10^{-7}$. For further analysis of the model, we choose
@@
$\ddot{\varphi}+3H\dot{\varphi}+U_{eff}^{(tree)\,\prime}=0,$ (3.17)
$H^{2}=\frac{1}{3M_{P}^{2}}\Bigl{(}\frac{1}{2}\dot{\varphi}^{2}+U_{eff}^{(tree)}(\varphi)\Bigr{)}$ (3.18)
@@
$3H\dot{\varphi}\simeq U_{eff}^{(tree)\,\prime}(\varphi),$ (3.19)
$H^{2}\simeq\frac{1}{3M_{P}^{2}}U_{eff}^{(tree)}(\varphi),$ (3.20)
@@
$\epsilon=\frac{M_{P}^{2}}{2}\Biggl{(}\frac{U_{eff}^{(tree)\,\prime}}{U_{eff}^{(tree)}}\Biggr{)}^{2}\ll 1,\qquad|\eta|=M_{P}^{2}\Biggl{|}\frac{U_{eff}^{(tree)\,\prime\prime}}{U_{eff}}\Biggr{|}\ll 1$ (3.21)
@@
At this stage of the study, our main goal is to understand how Higgs inflation with a small model parameter $\xi$ of nonminimal coupling becomes possible in the TMSM. First of all, choosing the value of $\xi=\frac{1}{6}$ we must find the value of the primordial parameter $\lambda$ in such a way that the model predictions agree with the CMB data. In light of what was said at the very end of the previous subsection, the predictions for the values of all perturbation parameters in the model under study must coincide with the corresponding predictions of the T-model. Therefore, to find $\lambda$, it is sufficient, for example, to consider the prediction for scalar perturbations $\mathcal{A}_{s}$. Using the standard method for estimating the duration of inflation, measured by the number of e-folds $N$, we obtain
$\epsilon\approx\frac{1}{8\xi N^{2}}$ (3.22)
@@
With the observational data bound [20], [21] $\mathcal{A}_{s}\approx 2.1\cdot 10^{-9}$ we get the relation
@@
Now, staying within the framework of a preliminary study, we will take the final step to demonstrate how the TMSM model proposed in this paper allows us to implement the Higgs inflation model with $\xi=\frac{1}{6}$ while satisfying the observed cosmological data due to $\lambda\sim 10^{-11}$, on the one hand, and to satisfy the requirement of $\lambda\sim 0.1$ near vacuum imposed by the SM, on the other hand. This is possible due to the arbitrariness of the primordial parameter $\lambda$ in TMSM, as discussed at the end of section 2.3. Indeed, as can be seen from eqs.(2.30) and (2.31), to obtain the desired VEV it suffices to choose the model parameter $b_{k}$ such that the smallness of $b_{k}-\zeta_{v}=b_{k}-1$ compensates for the smallness of $\lambda$. But as already noted in footnote 5, in ref. [22] it will be shown that the VEV of the Higgs field, which should be compared with the VEV in the SM, differs from the value of $\sigma$ given by eq.(2.31). However, in this paper, the exact value of $b_{k}-1$ does not play a significant role. Therefore, here it is sufficient to restrict ourselves to the order of magnitude of the result obtained in ref. [22]: $b_{k}-\zeta_{v}\sim 10^{-5}$. Thus, our goal is achieved by simply choosing a value for the parameter $b_{k}$ very close to $\zeta_{v}=1$. Turning to how the primordial action (2.1), (2.2) is structured, we understand that such a choice seems quite natural. Indeed, the extremely small deviation of $b_{k}$ from unity can be interpreted as a quantum effect of a more fundamental theory.

### v5

@@ 2.3 Vacuum of the Higgs field in the context of cosmology
The dependence of $U_{eff}^{(tree)}(\phi,\zeta;{\mathcal{M}})$, eq.(2.18), on $\zeta$, which, in turn, changes during the entire cosmological evolution (from inflation to transition to vacuum), plays a decisive role in all the physical results of the theory. Below, the TMT-effective potential will be obtained using eqs.(2.18) and (2.19). From now on we will assume that $V_{1}<0$, $V_{2}<0$ and $(electroweak\,scale)^{4}\lll|V_{1}|\sim|V_{2}|<M_{P}^{4}$ and we will see that by choosing the integration constant $\mathcal{M}$ one can ensure that the TMT-effective potential has a minimum at some value $\phi|_{min}=\sigma\lll M_{P}$. One can treat $\sigma$ as a vacuum expectation value (VEV) of $\phi$ obtained in the context of the studied cosmological model. The corresponding value $\zeta_{v}$ in the vacuum is obtained by making use of the constraint (2.19)
@@
Then eq.(2.21) reduces to an equation with a canonically normalized scalar field
@@
Here we neglected $\frac{\xi\sigma^{2}}{M_{P}^{2}}\lll 1$ in $\Omega$ and used eq.(2.10) rewritten in the Einstein frame. One should pay special attention that in order to obtain equation (2.23) in canonical form, it was necessary to devide eq.(2.21), considered near the vacuum, by a constant $\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}$. In secs.2.4 and 6.1 we will return to the discussion of the source of this question in more general case and will see that it is directly related to the K-essence type structure of the theory, which is responsible for the important results of the paper.
@@
where we introduced the notation $\Lambda({\mathcal{M}})$ to emphasize that with an appropriate choice of $\mathcal{M}$}, models with a very wide range of values of the cosmological constant can be realized. The natural constraint on $\Lambda$ is $0\leq\Lambda\lesssim(electroweak\,scale)^{4}$ and it is achieved if ${\mathcal{M}}^{2}\approx 4V_{1}V_{2}$. With the choice of the integration constant [^5] By choosing $\delta$ appropriately, one can ensure that $\Lambda=0$. This will exploit in section 7.3. However, this is nothing more than fine tuning, that is, in this way we do not solve the CC problem.
@@
$\sigma^{2}=\frac{\zeta_{v}-b_{p}}{b_{p}+\zeta_{v}}\cdot\frac{m^{2}}{\lambda}\left[1+\mathcal{O}\left(\frac{\lambda\sigma^{4}}{|V_{2}|}\right)\right].$ (2.30)
@@
and the notations [^6] We use the subscript $\sigma$ here to explicitly indicate that we are dealing with vacuum values of the TMT-effective parameters. Indeed, one might expect that, as follows from (2.21) and the constraint (2.19), $\phi$-dependent TMT-effective parameters $\lambda_{eff}$ and $m^{2}_{eff}$ appear as $\phi$ moves away from the vacuum. However, this is only partly true, namely for $\sigma\ll\phi\lesssim M_{P}$. As we will see in sections 2.4 and 4, substituting $\zeta$ from the constraint (2.19) into eq.(2.18) makes the TMT-effective potential essentially non-polynomial for $\phi\gtrsim M_{P}$, which does not allow us to use the terms TMT-effective parameters $\lambda_{eff}$ and $m^{2}_{eff}$ in the general case.
@@
are introduced for the values of quartic selfinteraction and squared mass parameters near the vacuum $\sigma$. Note that up to corrections of the order $\sim\mathcal{O}\left(\frac{\lambda\sigma^{4}}{|V_{2}|}\right)$
$\sigma^{2}=\frac{m_{\sigma}^{2}}{\lambda_{\sigma}}.$ (2.33)
If we restrict ourselves to studying the TMSM near vacuum, then the parameter $m$ will remain undefined, as well as the mass parameter in the GWS theory. The presence in the relations (2.32) of factors depending on the model parameters $\zeta_{v}=\sqrt{\frac{V_{1}}{V_{2}}}$, $b_{k}$ and $b_{p}$ means that there is a very wide range of these parameters in which the required VEV SM can be ensured [^7] However, in the model studied in this paper, we completely excluded gauge bosons and fermions and, accordingly, completely ignored the need to ensure gauge invariance. A suitably extended model that modifies the above results and allows adequate comparison with the GWS theory is studied in ref. [22], where it is shown that the VEV of the Higgs field, which should be compared with the VEV in the GWS theory, is not the value of $\sigma$ given by eq.(2.33), but the value of $\sqrt{\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}}\sigma$ (see also section 7.2). As will be shown in ref.[22], to demonstrate this it is necessary to take into account the differences in the description of TMSM in the so-called cosmological frame (CF) and in the local particle physics frame (LPPF). From this point of view, in studying cosmology in this paper, we limit ourselves to work in the CF, where the kinetic term in the TMT-effective Lagrangian is non-canonical, see below, eq.(2.41). The fundamentally new effect of the model proposed in this paper is that, thanks to these three model parameters, there is also freedom in the choice of the parameter $\lambda$. As we will see in section 3, this allows the implementation of the Higgs inflation model with a small constant $\xi$ of the non-minimal Higgs coupling to the scalar curvature.
@@ 2.4 General view of the TMSM
In section 2.2, TMT-effective equations in the Einstein frame were obtained. To understand the main features of TMSM at energies above the electroweak scale, it is necessary to return to these equations, taking into account the results of section 2.3 relating to what happens near the Higgs vacuum in the CF (see footnote 6). From the requirement that the TMT-effective cosmological constant be negligibly small, we obtained that, neglecting insignificant corrections, the integration constant is $\mathcal{M}=2\sqrt{V_{1}V_{2}}$, eq.(2.26). From now on, we will use the following parametrization for $V_{2}$
@@
Then the constraint (2.19) determines $\zeta=\zeta(\phi,X_{\phi})$ as follows
$\zeta=\frac{2q^{4}\zeta_{v}(1+\zeta_{v})-(2b_{p}-1)\frac{\lambda}{4}\frac{\phi^{4}}{M_{P}^{4}}-(2b_{p}+1)\frac{m^{2}}{2M_{P}^{2}}\frac{\phi^{2}}{M_{P}^{2}}+(1+b_{k})\Omega\frac{X_{\phi}}{M_{P}^{4}}}{2q^{4}(1+\zeta_{v})+\frac{\lambda}{4}\frac{\phi^{4}}{M_{P}^{4}}-\frac{m^{2}}{2M_{P}^{2}}\frac{\phi^{2}}{M_{P}^{2}}-(1+b_{k})\Omega\frac{X_{\phi}}{M_{P}^{4}}}.$ (2.35)
Substituting $\zeta(\phi,X_{\phi})$ into the $\zeta$-dependent form of $T_{\mu\nu}^{(eff)}$ in eqs.(2.17) and (2.18) with subsequent simple but rather lengthy algebraic calculations, the expression for the TMT-effective energy-momentum tensor reduces to
@@ 3.1 About the shape of the TMT-effective potential and other features of the model
The model contains 6 parameters $V_{1}$, $V_{2}$, $b_{p}$, $b_{k}$, $\lambda$ and $m$. Instead of $V_{2}<0$ it is convenient to use parameter $q^{4}\leq 1$ defined by the relation (2.34). Instead of $V_{1}<0$ we use $\zeta_{v}$ defined by eq.(2.27). It is natural to assume that $V_{1}$ and $V_{2}$ have the same order. Further in the paper, for numerical estimates and plotting graphs, we will need to select the values of $V_{1}$ and $V_{2}$. For simplicity, we choose
$V_{1}=V_{2}\quad\Longleftrightarrow\quad\zeta_{v}\equiv\sqrt{\frac{V_{1}}{V_{2}}}=1.$ (3.1)
Bearing in mind that the factor $\frac{V_{1}}{V_{2}}\sim\mathcal{O}(1)$ can lead to minor changes in the values of the calculated parameters, we can state that this choice is made without loss of generality. As a specific example of small $\xi$ from now on, we will always use $\xi=\frac{1}{6}$. We also choose
@@
However, as can be seen from eq.(2.41), the redefinition (3.3) absorbs only the factor $\Omega^{-1}$ in front of the first term, but does not provide canonical normalization of the field $\varphi$. The reason is that the TMT-effective action (2.40), (2.41) has the structure of K-essence type models, which, in turn, is due to the presence in the primordial TMSM action of two volume elements $dV_{g}$ and $dV_{\Upsilon}$ . Thus, we are limited to only the redefinition (3.3). The TMT-effective energy-momentum tensor given by eq.(2.36) now takes the following form
$T_{\mu\nu}^{(eff)}=K_{1}(\varphi)\cdot\big(\varphi_{,\mu}\varphi_{,\nu}-\tilde{g}_{\mu\nu}X_{\varphi}\big)-{\tilde{K}}_{2}(\varphi)\cdot\frac{X_{\varphi}}{M_{P}^{4}}\Big(\varphi_{,\mu}\varphi_{,\nu}-\frac{1}{2}\tilde{g}_{\mu\nu}X_{\varphi}\Big)+\tilde{g}_{\mu\nu}U_{eff}^{(tree)}(\varphi)$ (3.4)
and we obtain the TMT-effective action in terms of $\varphi$
@@
The last expression for ${\tilde{K}}_{2}(\varphi)$ takes into account that $X_{\phi}=\cosh^{2}z\cdot X_{\varphi}$.
@@
Figure 2: The curve labeled 1 is a plot of the TMT-effective potential $\frac{U_{eff}^{(tree)}(\varphi)}{M_{P}^{4}}\cdot 10^{10}$, defined by eqs.(3.7) and (3.8). It has one plateau due to the choice of $q^{4}=3\cdot 10^{-10}$, i.e. $|V_{1}|=|V_{2}|\approx(10^{16}GeV)^{4}$, while other model parameters are the same as in Fig. 1. The curves labeled 2 and 3 are plots of the scalar function $\zeta(\varphi)$ defined by the constraint (3.1), where the terms $\propto\frac{X_{\varphi}}{M_{P}^{4}}$ are assumed to be negligible compared to the other terms. Curve 2 is the plot of $\zeta(\varphi)$ if the model parameter $b_{p}=0.55$ is chosen, and the plot intersects $\zeta=0$ at $\varphi=4.25M_{P}$. Curve 3 is the plot of $\zeta(\varphi)$ if the model parameter $b_{p}=0.5(1+10^{-8})$ is chosen. In this case, curve 3 intersects $\zeta=0$ at $\varphi\approx 14M_{P}$.
@@
    Let us now analyze in more detail the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$ with one plateau. With the choice described by eqs.(3.1) and (3.2), and using the parameter values $\lambda=2.3\cdot 10^{-11}$,   $\xi=\frac{1}{6}$ tested in sec.3.2, and $m=0.7GeV$, one can see that the TMT-effective potential with one plateau is reached [^9] For $\varphi>\frac{M_{P}}{\sqrt{\xi}}=\sqrt{6}M_{P}$, we find that $F(\varphi)\approx\frac{1}{2}$ with a relative error $\frac{|F(\varphi)-\frac{1}{2}|}{1/2}<2.2\cdot 10^{-3}$. if $q^{4}\lesssim 10^{-7}$. For further analysis of the model, we choose
@@
$\ddot{\varphi}+3\frac{\dot{a}}{a}\dot{\varphi}+U_{eff}^{(tree)\,\prime}=0,$ (3.17)
$\frac{\dot{a}^{2}}{a^{2}}=\frac{1}{3M_{P}^{2}}\Bigl(\frac{1}{2}\dot{\varphi}^{2}+U_{eff}^{(tree)}(\varphi)\Bigr)$ (3.18)
@@
$3\frac{\dot{a}^{2}}{a^{2}}\dot{\varphi}\simeq U_{eff}^{(tree)\,\prime}(\varphi),$ (3.19)
$\frac{\dot{a}^{2}}{a^{2}}\simeq\frac{1}{3M_{P}^{2}}U_{eff}^{(tree)}(\varphi),$ (3.20)
@@
$\epsilon=\frac{M_{P}^{2}}{2}\Biggl(\frac{U_{eff}^{(tree)\,\prime}}{U_{eff}^{(tree)}}\Biggr)^{2}\ll 1,\qquad|\eta|=M_{P}^{2}\Biggl|\frac{U_{eff}^{(tree)\,\prime\prime}}{U_{eff}^{(tree)}}\Biggr|\ll 1$ (3.21)
@@
At this stage of the study, our main goal is to understand how Higgs inflation with a small model parameter $\xi$ of nonminimal coupling becomes possible in the TMSM. First of all, choosing the value of $\xi=\frac{1}{6}$ we must find the value of the primordial parameter $\lambda$ in the TMT-effective potential (3.16) in such a way that the model predictions agree with the CMB data. In light of what was said at the very end of the previous subsection, the predictions for the values of all perturbation parameters in the model under study must coincide with the corresponding predictions of the T-model. Therefore, to find $\lambda$, it is sufficient, for example, to consider the prediction for scalar perturbations $\mathcal{A}_{s}$. Using the standard method for estimating the duration of inflation, measured by the number of e-folds $N$, we obtain
$\epsilon\approx\frac{1}{8\xi N^{2}}\qquad\text{and}\qquad\eta\approx-\frac{1}{N}.$ (3.22)
@@
With the observational data bound [20] $\mathcal{A}_{s}\approx 2.1\cdot 10^{-9}$ we get the relation
@@
With $\xi=\frac{1}{6}$, for the scalar spectral index $n_{s}=1-6\epsilon+2\eta$ and the tensor-to-scalar ratio $r=16\epsilon$ we find for $N=60$
$n_{s}=1-\frac{2}{N}-\frac{3}{4\xi N^{2}}\approx 0.965\qquad\text{and}\qquad r=\frac{2}{\xi N^{2}}\approx 0.003,$ (3.26)
which are in good agreement with the observational results of the Planck 2018 survey [20] and the BICEP/Keck measurement [21].
Now, staying within the framework of a preliminary study, we will take the final step to demonstrate how the TMSM model proposed in this paper allows us to implement the Higgs inflation model with $\xi=\frac{1}{6}$ while satisfying the observed cosmological data due to $\lambda\sim 10^{-11}$, on the one hand, and to satisfy the requirement of $\lambda\sim 0.1$ near vacuum imposed by the SM, on the other hand. This is possible due to the arbitrariness of the primordial parameter $\lambda$ in TMSM, as discussed at the end of section 2.3. Indeed, as can be seen from eqs.(2.32) and (2.33), to obtain the desired VEV it suffices to choose the model parameter $b_{k}$ such that the smallness of $b_{k}-\zeta_{v}=b_{k}-1$ compensates for the smallness of $\lambda$. But as already noted in footnote 5, in ref. [22] it will be shown that the VEV of the Higgs field, which should be compared with the VEV in the SM, differs from the value of $\sigma$ given by eq.(2.33). However, in this paper, the exact value of $b_{k}-1$ does not play a significant role. Therefore, here it is sufficient to restrict ourselves to the order of magnitude of the result obtained in ref. [22]: $b_{k}-\zeta_{v}\sim 10^{-5}$. Thus, our goal is achieved by simply choosing a value for the parameter $b_{k}$ very close to $\zeta_{v}=1$. Turning to how the primordial action (2.1), (2.3) is structured, we understand that such a choice seems quite natural. Indeed, if we recall that the volume measure with which the kinetic term enters the primordiial action (2.3), $(b_{k}\sqrt{-g}-\Upsilon)d^{4}x=\sqrt{-g}(b_{k}-\zeta)d^{4}x$, the extremely small deviation of $b_{k}$ from the vacuum value of $\zeta$, i.e. $\zeta_{v}$, can be interpreted as the effect of the violation of some unknown symmetry of a more fundamental theory.